OpenStudy (anonymous):

Please help! Write the equation 2x+7y-4=0 in normal form.

OpenStudy (anonymous):

@SolomonZelman Can you please help me?

TheSmartOne (thesmartone):

What do you mean by normal form?

TheSmartOne (thesmartone):

Standard from??

OpenStudy (anonymous):

like this: A. \[A, \frac{ 2x }{ \sqrt69 } + \frac{ 7y }{\sqrt69 } - \frac{ 4 }{ \sqrt69 } =0\]

OpenStudy (anonymous):

\[B. \frac{ 2x }{ \sqrt53 } + \frac{ 7y }{\sqrt53 } - \frac{ 4 }{ \sqrt53 } =0\]

OpenStudy (anonymous):

\[C. \frac{ 2x }{ \sqrt13 } + \frac{ 7y }{\sqrt13 } - \frac{ 4 }{ \sqrt13 } =0\]

OpenStudy (anonymous):

\[D. \frac{ 2x }{ \sqrt9 } + \frac{ 7y }{\sqrt9 } - \frac{ 4 }{ \sqrt9 } =0\]

TheSmartOne (thesmartone):

I have no clue.. Sorry

OpenStudy (anonymous):

It fine. I got it now :)

OpenStudy (anonymous):

Thanks for trying to help!

TheSmartOne (thesmartone):

No problem :D